# LLM-Generated Heuristics for AI Planning: Do We Even Need Domain-Independence Anymore?

### v3

@@ Introduction
LLM-generated heuristics are not always compilable or computationally efficient, occasionally causing the search to exceed memory or time constraints (10 minutes and 8 GB of memory). To address this, we employ two fallback strategies. In FirstCompilation (fc), we repeatedly query the LLM until a compilable heuristic is obtained and run GBFS with it until a solution is found or resources are exhausted, 600 sec (search + API calls). In TimeSlicedRestarts5 (tsr5), we allocate fixed 100 sec time slices to successive heuristics; if a run fails or exceeds memory, we restart with a new heuristic. This continues until a solution is found or the time limit is reached. All API call durations, typically 24–27 seconds, are included in the time budget.
@@
We evaluated four LLMs for this task: OpenAI’s GPT-4.1 and GPT-4.1-mini, and Anthropic’s Claude Sonnet 3.5 and Claude Haiku 3.5. Results are presented in Table 1. As is standard in AI planning, we report coverage rather than averages, since state space size typically grows exponentially and averaging would be misleading. The head-to-head time for solution comparison of tsr5 with GPT-4.1 vs. the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$ can be seen in Figure 2.
Ablation Study
Qry 1 2 3 4 5 6 7 8
4.1
U 199 265 281 281 282 282 282 282
R 162 215 247 258 265 - - -
B - 239 - 275 - 283 - 284
mini
U 194 203 207 228 228 247 247 247
R 121 170 194 210 224 - - -
B - 156 - 205 - 226 - 238
Table 2: Ablation: For GPT-4.1 and GPT-4.1-mini, we compare three configurations: using additional unrefined heuristics only (U), using refined heuristics only (R), and combining both unrefined and refined heuristics (B). The configuration with refined heuristics 6–8 was not run due to time constraints (see Supplementary Materials). We also note that the unrefined version did not improved for up to 10 iterations.
@@
We conducted an ablation study to evaluate the contribution of two components in our heuristic generation pipeline: (i) the heuristic refinement phase and (ii) the preceding strategizing step. The study aimed to determine their impact on heuristic quality and planning performance, as well as the variance in our best configuration. Due to space constraints, results are reported for GPT-4.1; additional configurations are provided in the Supplementary Materials.
Impact of Refinement
The third phase of the pipeline refines the heuristic using instance-specific information from a JSON file. This is achieved by prompting the LLM with both the unrefined heuristic and problem-specific data. We compare the refined heuristics against their unrefined versions.
Empirically, refinement leads to a slight increase in overall coverage (Table 2). However, closer inspection reveals that this improvement stems primarily from increased variance across runs: by combining heuristics generated at different phases (unrefined and refined), the likelihood of obtaining at least one successful variant increases. In other words, the benefit is less a result of systematic improvement and more a consequence of broader sampling.
Impact of Strategizing
We performed an optional first “strategizing” phase, where the LLM is prompted to verbally reason about how to create an effective heuristic for the domain. Reasons for inclusion of this step are described in Methodology section. Surprisingly, in practice we find that despite it’s Chain of Thought step-by-step nature, it actually hurts performance (Table 3).
Letting the model generate code from the start yielded better coverage and required fewer tokens. We observe no consistent improvements in heuristic structure or robustness from allowing strategizing. We attribute this partly to it’s performance on Sailing where it solved 19 instances.
Strat. No Strat.
IPC-23 domains only 247 253
Total 282 288
Avg. input tokens 56.0k 41.3k
Avg. output tokens 8.8k 6.8k
Table 3: Comparison of strategizing vs. no strategizing in terms of token usage and instance counts for TSR5 with GPT-4.1. The average tokens are reported per instance.
Average GPT Claude
TSR 4.1 4.1-mini 4o 4o-mini Sonnet 3.5 Haiku 3.5
Input tokens 56074 71202 66124 72407 76064 90115
Completion tokens 8880 11311 7055 7163 8105 8971
Compilation failure 0.107 0.368 0.122 0.339 0.089 0.373
OutOfMemory rate 0.304 0.241 0.471 0.369 0.448 0.312
Table 4: Token usage and compilation failure rates. We report averages for input tokens and completion tokens per heuristic, and the percentage of compilation failures.
Figure 3: Variance per domain for the tsr5 GPT-4.1 unrefined configuration over three independent runs, each limited to 600 seconds and 8 GB of memory per instance. Total coverages were 282, 282, and 276, indicating stable performance. All runs outperform the domain-independent SOTA. For comparison, we also include results from a run without the strategizing phase.
Summary
The ablation results suggest that refinement has minor positive effects, primarily due to increasing the diversity of heuristic candidates. The strategizing step, however, does not yield measurable benefit. These findings support the use of simple multi-sampling strategies to increase coverage, while questioning the need for more elaborate prompting constructs in heuristic generation.
@@ Discussion
As shown in Table 1, LLM-based heuristics achieve state-of-the-art performance across the evaluated domains. They are particularly effective in challenging benchmarks such as Sailing and Zenotravel, which are known to be difficult for domain-independent planners. The best-performing LLM-based heuristics were produced by GPT-4.1, with Sonnet 3.5 yielding comparable results. Nonetheless, certain domains remain challenging for LLMs, likely due to their sensitivity to non-descriptive variable names and opaque goal specifications, as seen in Pathways and Settlers.
@@
Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (Stojkovic et al. 2024), energy efficiency is beyond the scope of this work. Nonetheless, our method is amenable to efficient deployment: the first two phases are domain-specific rather than instance-specific and can be executed once per domain, substantially reducing energy and cost. On average, solving an input using TSR5 with GPT-4.1 requires 56k input tokens and 8.8k output tokens (Table 4), corresponding to a pre-caching cost of approximately 0.18 USD.

### v4

@@ Introduction
In SelfPortfolio-N (sp-n), we allocate fixed $Time/N$s slices to N heuristics; if a run fails or exceeds memory, we restart with generating a new heuristic. This continues until a solution is found or all slices failed. All API call durations, typically 10–45 seconds depending on the model, are included in the time budget.
@@
We evaluated The following LLMs and settings: OpenAI’s GPT-4.1 and GPT-4.1-mini, their reasoning models GPT-5.1 and GPT-5-mini (under both a ”low” and ”high” reasoning effort setting), and Anthropic’s Claude Sonnet 4.5 and Claude Haiku 4.5. Results are presented in Table 1. For practicality, we limit the maximum number of heuristics generated for fc and us to 10, which if passed the instance is declared failed. As is standard in AI planning, we report coverage rather than averages, since state space size typically grows exponentially and averaging would be misleading. As a qualitative measure of the generated heuristics we report the head-to-head time and expanded nodes for solution comparison of sp-10 with GPT-4.1 vs. the best performing $h^{\text{add}}_{\langle\text{B, QB}\rangle}$, which can be seen in Figure 5.Comparisons with additional models can be seen in Appendix.
@@
We conducted ablation studies to evaluate the contribution of three components in our heuristic generation pipeline:(i) the provision of the specific instance (ii) the model and fallback choice (iii) the reasoning effort requested from reasoning models. The study aimed to determine their impact on heuristic quality, generation time, and planning performance, as well as the variance in our best configuration.
Impact of Instance-Specificity
We measure the effect of providing the instance JSON to the model before generation, containing the initial state and the goal. This comes at the steep cost of requires generating additional heuristics per instance, as opposed to domain-specific-instance-general which can be generated once then ran on all future instances. Intuitively, IS may have the following benefits:
  1. 1.
  2. 2.
    It tells the model how significant each part of the problem is. Consider a simple instance of Pacman where the ghosts are trapped. The model can avoid spending generation- and run-time on a complex subroutine to quantify the risk from the ghosts.
  3. 3.
    It tells the model the scale of the problem which can provide guidance on efficiency constraints. We hypothesize that there exists a tradeoff between the precision of the heuristic estimates and the computation time, and further that for larger problems it can be worthwhile for the heuristic to sacrifice precision for efficiency. For example, on TPP, Anthropic’s models implement path-finding algorithms like Dijkstra, Prim and Floyd-Warshall, about half the time. Although this is a good basis for a “smart” heuristic, it causes a timeout within a few thousand expansions for even medium problems. Meanwhile even Manhattan Distance heuristics achieves 16/20.
  4. 4.
    At times, a solution state is obvious from the initial state (e.g, FO-Counters). In these cases showing the model the instance details allows it to write a heuristic to guide efficiently toward that solution rather than abstractly guiding toward all solutions.
Empirically, IS leads to a decrease in overall coverage (Table 2) with most domains varying slightly but a few increasing or decreasing significantly. IS did not cause heuristic generation time to increase [^9] Interestingly, despite adding tokens and theoretically adding complexity which can increase output tokens, adding the instance slightly decreased generation time. (P=2e-5<<0.05). We do not have an explanation for this fact.. IS moderately increased compilation errors (see Table  3), and if this effect is normalized against, IS-heuristics have a slightly higher solving rate. We hypothesize that GPT-4.1 is insufficiently strong to make use of this info and that for a stronger model it may be worth revisiting.
FC US SP-10
Domain $D$ IS $D$ IS $D$ IS
Block Grouping (20) 16 9 19 10 19 10
Counters (20) 10 8 10 11 10 11
Delivery (20) 16 10 20 18 20 18
Drone (20) 15 15 16 16 16 16
Expedition (20) 3 1 3 4 3 4
Farming (20) 18 19 19 20 19 20
FO-Counters (20) 4 8 4 14 4 14
FO-Farming (20) 20 18 20 20 20 20
FO-Sailing (20) 20 10 20 19 20 20
Hydropower (20) 8 7 12 9 12 9
Market Trader (20) 20 1 20 3 20 3
Pathways (20) 0 1 1 1 1 1
Plant Watering (20) 9 12 20 18 20 18
Rover (20) 3 2 4 3 4 3
Sailing (20) 11 14 20 20 20 20
Settlers (20) 1 2 1 3 3 3
TPP (20) 13 11 16 13 16 15
Zenotravel (20) 20 14 20 20 20 20
$\sum$ (360) 207 162 245 222 247 225
Twin Prime (20) 18 20 18 20 18 20
Pacman (20) 9 13 9 13 16 14
$\sum$ (400) 234 195 272 255 281 259
Table 2: Coverage results for GPT-4.1 with domain-dependence only ($D$), our standard configuration, against GPT-4.1 with Instance-Specificity (IS), reporting each on fc, us, and their best configuration, SP-10. For most domains GPT-4.1 fails to leverage the additional information, however we highlight FO-Counters and Twin Prime as domains where it lead to the most improvement.
Impact of model and fallback choice
We tested 6 models: GPT-4.1, GPT-4.1-mini, GPT-5.1, GPT-5-mini, Claude Sonnet 4.5, and Claude Haiku 4.5, each with each of our fallback options including fc, us, sp-5, and sp-10. For the full table see Appendix. Our observations are as follows:
  1. 1.
    For non-reasoning models, under about every configuration, a model’s larger variant is preferred. Since they are well-capable of generating and running many heuristics within the allotted time, SP10SP-10 is preferred.
  2. 2.
    Reasoning models are more competitive between sizes. They take longer to generate and $US$ performs best, allowing them the time they need.
  3. 3.
    Anecdotally, when reasoning models fail it can often be because they over-complicate the problem, taking longer to generate, and whether they end up with a better heuristic or an unusable one is a coin toss. That said, since we can re-sample this risk may be worth taking if the samplings are sufficiently independent.
Impact of Reasoning effort
We tested our method on the same model, GPT-5.1 with two configurations: reasoning_effort=’low’ and reasoning_effort=’high’ (for brevity denoted ’GPT-5.1’ and ’GPT-5.1=high’). We found that although increasing reasoning_effortreasoning\_effort slightly increased heuristic efficiency and quality, and better-than-halved compilation error rate, when we account for generation time the reduced number of heuristics we generate in time does not make this trade worthwhile. With better engineering one could generate many completions simultaneously, which may turn this trade worthwhile. We leave this direction for future work. For further graphs see appendix.
GPT Claude
Metric 4.1 4.1 (IS) 4.1 mini 5.1 5 mini 5.1 high 5 m-h Sonnet 4.5 Haiku 4.5
Input tok. 3725 8102 3725 3724 3724 3724 3724 4761 4761
Output tok. 1510 1695 1245 3253 2298 18075 14485 2122 1490
Reasoning tok. [^10] For GPT, subset of output tokens. For Claude, they are not provided by the API - - - 1282 589 15510 11613 - -
Cost (USD) 0.02 0.03 0.003 0.037 0.006 0.185 0.03 0.046 0.012
Generation (sec.) 35.8 27.5 19.4 44.9 44.5 235.1 202.3 33.6 12.5
Run (sec.) 24.7 22.1 19.2 24.2 18.0 21.4 36.7 43.4 22.2
Compl. error % 0.300 0.430 0.215 0.150 0.405 0.070 0.205 0.195 0.165
OOM rate % 0.304 0.241 0.388 0.382 0.280 0.340 0.288 0.391 0.451
Runtime error % 0.027 0.020 0.035 0.032 0.028 0.035 0.032 0.016 0.032
Timeout % 0.015 0.014 0.004 0.007 0.009 0.007 0.031 0.032 0.002
Success % 0.355 0.295 0.358 0.430 0.278 0.548 0.444 0.365 0.351
Table 3: Average token usage, cost, latency, and execution outcomes for different GPT and Claude models, independent of downstream applications. Models marked with are reasoning models. Excepting IS, input tokens across models depend only on the tokenizer. Costs do not account for token caching discounts.
Figure 4: As a variance analysis we generated 40 heuristics with GPT-4.1 in each domain, and performed a Monte-Carlo simulation of our algorithm for 1000 iterations of sampling heuristic order, taking the coverage per-domain and overall each time. The boxes represent the first and third quartiles, the thick line the median, and the whiskers the 1st and 99th percentiles. Although limited, the analysis shows the method to be highly consistent for most domains.
The ablation results suggest that non-reasoning models don’t significantly benefit from IS. The best setting of our method is to use a non-reasoning model to generate non-IS heuristics and run them as SelfPortfolio-10 (276 with GPT-4.1). If time is allowed ahead of receiving the instances, the best setting becomes instead to generate ahead of time as many heuristics as possible and run SP-N or US on them (The two fallbacks are approximately equivalent when used on pregenrated heuristics due to near-zero timeout rate).
@@ Discussion
As shown in Table 1, LLM-based heuristics achieve state-of-the-art performance across the evaluated domains. They are particularly effective in challenging benchmarks such as (FO-) Sailing and Zenotravel, which are known to be difficult for domain-independent planners. The best-performing LLM-based heuristics were produced by GPT-5.1 and it’s mini variant with high reasoning effort, however when accounting for generation time and cost, GPT-4.1 emerges as a clear winner. Nonetheless, certain domains remain challenging for LLMs, likely due to their sensitivity to non-descriptive variable names and opaque goal specifications, as seen in Pathways and Settlers.
@@
Costs and Efficiency While concerns about the resource efficiency of LLMs have been raised (stojkovic2024towards), energy efficiency is beyond the scope of this work. Nonetheless, our method is amenable to efficient deployment: generating heuristics can be done ahead of time one per domain then reused on every instance, substantially reducing energy and cost even if a powerful model is used. Additionally, the resulting heuristics tend to be very efficient to run, and succeed or fail much quicker than domain-independent heuristics (see Appendix.)